OpenStudy (seattle12345):

can someone help me with this: Springfield and Futureland are 685 miles apart. At noon, Melanie leaves Springfield traveling toward Futureland at 50 miles per hour. 0.5 hours later, Sacheen leaves Futureland heading for Springfield at 60 miles per hour. How long before they will pass each other on the road?

hero (hero):

Another systems of equations

hero (hero):

Are you there?

hero (hero):

Yahya?

OpenStudy (seattle12345):

Hi, yea sorry. I'm confused on how you came up with 685 = 685 50t = 60t - 30 though

OpenStudy (seattle12345):

oh I see. I just so used to seeing y=x+B lol. I ended up with 6 when I attemped to answer this. Its say the answer is 6.5 though.

OpenStudy (campbell_st):

its an an interesting answer t = 3, who actually has been travelling for 3 hours melanie or sacheen...? I'll leave you to figure it out

hero (hero):

M = 50t S = 60t - 30 D = 685 t = time it takes for Melanie and Sacheen to meet 50t + (60t - 30) = 685 50t + 60t - 30 = 685 110t - 30 = 685 110t = 685 - 30 110t = 655 t = 655/110 t = 5.9545... t is approximately 6 hours

hero (hero):

I rushed through the question too fast without reading it properly the first time

OpenStudy (campbell_st):

thanks hero... makes more sense now

hero (hero):

M = 50t S = 60t - 30 D = 685 t = time it takes for Melanie and Sacheen to meet 50t + (60t - 30) = 685 50t + 60t - 30 = 685 110t - 30 = 685 110t = 685 + 30 110t = 715 t = 715/110 t = 6.5 t = 6.5 Actually I made yet another mistake. But t is correct now.

OpenStudy (seattle12345):

Hero... you are amazing lol!

hero (hero):

Nah, average at best considering the mistakes

OpenStudy (seattle12345):

@campbell_st sacheen has been traveling 30 min less than melanie wich is 6 hours and they pass each other at 6:30 p.m.

OpenStudy (seattle12345):

well you corrected them so above average now lol